When MOSCA was mounted it was found that the bias levels of the different CCDs had changed (again), but this time also the gain of one of the CCDs (W20) had changed by 20%. This was rather worrisome and a full linearity test was made. All the CCDs showed very good linearity and normal RON levels, so it was decided to leave the instrument on the telescope. Also, cards and cables where swapped between the four CCDs without any indication of a problem. No specific problems with the detectors were reported during the run.
A procedures to test the CCDs, and analyze and present the results on the web similar to that for ALFOSC was implemented. This was actually extremely useful when investigating the varying gain level mentioned above.